32 1/3 =2x

321/3=2x

Grade:8

3 Answers

Nishant Vora IIT Patna
8 years ago
Dear student
Pls find the solution

2 ^ (5/3) = 2x

x = 2 ^ (5/3 – 1)

Thanks
noogler
8 years ago
32=25
sub it, u get 25/3= 2x
                   x=25/3/2=25/3-1=22/3
OUTPASS
8 years ago
2 ^ (5/3) = 2xx = 2 ^ (5/3 – 1)